Question three

The results in Management examinations at ICU were as shown below:

63

65

67

68

69

70

71

72

74

75

x

If the mean of this data is 69

Calculate:

  1. The value of x (2 marks)
  2. Interquartile range (4 marks)
  3. The intelligence of student was measured and using the ranking of 5 as excellent, 4 Very good, 3 Good 2 Average and 1 below average

The table below shows distribution of intelligence with grades:

Grades

63

65

67

68

69

70

71

72

74

75

Intelligence

1

1

1

1

3

3

4

4

5

5

  1. Using Pearson’s product moment correlation coefficient determine the relation between grades and intelligence consider intelligence as independent variable (12 marks)
  2. Define statistics (2 marks)

    Question three

    The results in Management examinations at ICU were as shown below:

    63

    65

    67

    68

    69

    70

    71

    72

    74

    75

    x

    If the mean of this data is 69

    Calculate:

  3. The value of x (2 marks)
  4. Interquartile range (4 marks)
  5. The intelligence of student was measured and using the ranking of 5 as excellent, 4 Very good, 3 Good 2 Average and 1 below average
  6. The table below shows distribution of intelligence with grades:

    Grades

    63

    65

    67

    68

    69

    70

    71

    72

    74

    75

    Intelligence

    1

    1

    1

    1

    3

    3

    4

    4

    5

    5

  7. Using Pearson’s product moment correlation coefficient determine the relation between grades and intelligence consider intelligence as independent variable (12 marks)
  8. Define statistics (2 marks)
Answer #1

SOI?:) Value of x ! - Mean = sum of all observations Total no. of obseri 69 = 63+ 65 +67+68+69+ 70- +74+75+7 : 69 = 69472 11N = Ey qax = X = 699 69.4. - 23 - lo x= Grade Y = Intelligence.. x 491 229 ny y @ L !!ī ī. -ammus x= x - x y=Y-I 1 - 64 -1.8» Correlation wefficient rla,y) is VEz? 53.8 √130. 4 √25.6 53.8 (1):4192) (5,0596) = 0.931) Since r(wy) -0.9311 , the variceb